OpenStudy (anonymous):

Solve for x: 8^(5x-10)=(1/16)^(2x-7)

OpenStudy (anonymous):

Can you explain how you got to that?

OpenStudy (anonymous):

u can do this

OpenStudy (anonymous):

2^3^(5x-10) = (2^-4)^(2x-7) since 8 =2^3 and 1/16= 2^-4

OpenStudy (anonymous):

so u get 2^15x-30 = 2^-8x+28

OpenStudy (anonymous):

now ur bases are the same so u only have to equate the indices

OpenStudy (anonymous):

15x-30= -8x+28

OpenStudy (anonymous):

Ok.That makes sense. Thanks a lot.

OpenStudy (anonymous):

23x=58 x=2.52